OpenStudy (anonymous):

Implicit differentiation: y^2 = (x^2)/(xy-4). Thank you!

OpenStudy (anonymous):

ick

OpenStudy (anonymous):

\[y^2=\frac{x^2}{xy-4}\] right?

OpenStudy (anonymous):

yes! ick is right haha

OpenStudy (anonymous):

the best way to think of this is as \[f^2(x)=\frac{x^2}{xf(x)-4}\] you need the product rule, the quotient rule, and the chain rule

OpenStudy (anonymous):

the left hand sides is easy, the derivative of \(f^2(x)\) is \(2f(x)f'(x)\) which you write as \[2yy'\]

OpenStudy (anonymous):

the right hand side is a bear

OpenStudy (anonymous):

yeah, the right part is the... .. T__T

OpenStudy (nincompoop):

love your sarcasm, @satellite73

OpenStudy (anonymous):

ok but it is still doable the derivative of the numerators is \(2x\) and the derivative of the denominator is \[y+xy'\] by the product rule @nincompoop i'll take that as a compliment thank

OpenStudy (nincompoop):

laughing out loud

OpenStudy (anonymous):

is it fine if it change it to: \[(y ^{2})(xy-4)-x^2=0\]?

OpenStudy (anonymous):

you get via the quotient rule \[\frac{(xy-4)\times 2x-x^2(y+xy')}{(xy-4)^2}\]

OpenStudy (anonymous):

then comes a raft of algebra i guess that might help we can try it both ways

OpenStudy (anonymous):

actually that might be a good idea we get \[(y ^{2})(xy-4)-x^2=0\] or \[xy^3-4y^2-x^2=0\] lets see \[y^3-3xy^2y'-9yy'=2x=0\] hmm looks promising

OpenStudy (anonymous):

\[y^3-3xy^2y'-8yy'-2x=0\]

OpenStudy (anonymous):

then \[y^3-2x=3xy^2y'-8yy'\] \[y^3-2x=y'(3xy^2-8y)\] \[y'=\frac{x^3-2x}{3xy^2-8y}\] well that wasn't so bad was it hope it is right shall we check?

OpenStudy (anonymous):

hmm, looks good!!

OpenStudy (anonymous):

yeah too bad it is not the wolfram answer

OpenStudy (anonymous):

ah. if only this wasn't an even numbered question ..

OpenStudy (anonymous):

lets try again

OpenStudy (anonymous):

thanks man!! I super appreciate your effort haha.

OpenStudy (anonymous):

first of all i made a stupid algebra mistake

OpenStudy (anonymous):

if we go ahead with \[(y^2)(xy-4)-x^2=0 \]then we could try to use product rule twice?

OpenStudy (anonymous):

it will be a lot lot easier to multiply out first

OpenStudy (anonymous):

\[xy^3-4y^2-x^2=0\] as a first step then \[y^3-3xy^2y'-8yy'-2x=0\]

OpenStudy (anonymous):

then \[y^3-2x=3xy^2y'+8yy'\] \[y'=\frac{y^3-x}{3xy^2+8y}\] but that is still not right!

OpenStudy (anonymous):

wolf has a \(-8y\) in the denominator where i have it as plus but a totally different answer if you use hte quotient rule

OpenStudy (anonymous):

hm, does wolf talk into account implicit differentiation?

OpenStudy (anonymous):

i am going back to this \[2yy'=\frac{(xy-4)\times 2x-x^2(y+xy')}{(xy-4)^2}\]

OpenStudy (anonymous):

yes it does it will find both \(\frac{dy}{dx}\) and also \(\frac{dx}{dy}\)

OpenStudy (anonymous):

i did out question and got wolfram answer. 1 sec i will post picture

OpenStudy (anonymous):

\[2yy'=\frac{(xy-4)\times 2x-x^2(y+xy')}{(xy-4)^2}\] \[(xy-4)^2\times 2yy'=2x^2-8x-x^2y-x^3y'\]

OpenStudy (anonymous):

OH MY.

OpenStudy (anonymous):

\[2y(xy-4)^2y'+x^3y'=2x^2-8x-x^2y\] \[y'=\frac{2x^2-8x-x^2y}{2y(xy-4)^2}\] is what i get

OpenStudy (anonymous):

no another algebra mistake

OpenStudy (anonymous):

\[y'=\frac{2x^2-8x-x^2y}{2y(xy-4)^2+x^3}\]

OpenStudy (anonymous):

oh well, guess my algebra is not up to the task

OpenStudy (anonymous):

it was a great effort!! I reward you.

OpenStudy (anonymous):

actually i think it might be right

OpenStudy (anonymous):

hopefully you can see that picture

OpenStudy (anonymous):

yeah did you get that from wolf? i think they might be the same

OpenStudy (anonymous):

wow, haha. thank you!!

OpenStudy (anonymous):

i worked it out myself but it matches wolf answer

OpenStudy (anonymous):

i wish i could give 2 rewards.. this question was too crazy

OpenStudy (anonymous):

no worries. can you help me with my implicit differentiation question i have?

OpenStudy (anonymous):

then you win i can't find my mistake though...

OpenStudy (anonymous):

guess i need pencil and paper too, stop trying to write it all here

OpenStudy (anonymous):

@amyvincent12 what is your question?

OpenStudy (anonymous):

y=csc(xy) find y' using implicit differentiation

OpenStudy (anonymous):

\[y'=-\cos(xy)\cot(xy)(y+xy')\] is a start

OpenStudy (anonymous):

i have that

OpenStudy (anonymous):

ok i hope not, it is wrong i meant \[y'=-\csc(xy)\cot(xy)(y+xy')\]

OpenStudy (anonymous):

then multiply out and get \[y'=-y\csc(xy)\cos(xy)-xy'(\csc(xy)\cot(xy)\]

OpenStudy (anonymous):

sorry, didn't see typo. yes, this is what i have.

OpenStudy (anonymous):

put all the \(y'\) on the left and get \[y'+xy'(\csc(xy)\cot(xy)=-y\csc(xy)\cot(xy)\]

OpenStudy (anonymous):

factor out the \(y'\) to get \[y'(1+x\csc(xy)\cot(xy))=-y\csc(xy)\cot(xy)\]

OpenStudy (anonymous):

then divide and you are done

OpenStudy (anonymous):

look reasonable?

OpenStudy (anonymous):

i got an answer of \[y'=\frac{ -ycsc(xy)\cot(xy) }{1+xcsc(xy)\cot(xy)}\]

OpenStudy (anonymous):

but doesn't match wolfram

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

wolfram doesn't have denominator at all.

OpenStudy (anonymous):

well, there are lots of trig identities you could use maybe to change it let me check

OpenStudy (anonymous):

the only trig identities that my prof wants us to change is sinx/cosx =tanx and cosx/sinx = cotx so thats probably my final answer just not for wolf

OpenStudy (anonymous):

wolf gets the same answer i do

OpenStudy (anonymous):

okay awesome maybe i made typo in wolf… thanks!!!!!

OpenStudy (anonymous):

actually when i type it in to wolf it gives \[y'=y'(x) = -\frac{(y\cos(x y))}{\sin^2(x y)+x\cos(x y)}\] but when i got to "step by step" it gives my answer above hmmm

OpenStudy (anonymous):

probably some messing around with trig that get it, i don't see it

OpenStudy (anonymous):

probably replace the 1 in the denominator by \(\sin^2(xy)+\cos^2(xy)\) and then some algebra